OpenStudy (anonymous):

Simplify and write the expression with positive exponents only. b^-3/b^-5

OpenStudy (anonymous):

use the rule \[x ^{-a}=\frac{ 1 }{ x ^{a} }\]

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

wait this is different because it has the variables and fractions.. idk how to set it up with the formula :( can you help me again?

OpenStudy (anonymous):

\[\frac{ b ^{-3} }{ b ^{-5} }=\frac{ b ^{5} }{ b ^{3} }\]

OpenStudy (anonymous):

ohhhh okay thank you!

OpenStudy (anonymous):

do i just divide 5 and 2 after that?

OpenStudy (anonymous):

the rule for that \[\frac{ x ^{a} }{ x ^{b} }=x ^{a-b}\]

OpenStudy (anonymous):

oh nvm then i had a diff one written down

OpenStudy (anonymous):

but what do i do with the exponents of 5 and 2?

OpenStudy (anonymous):

\[b ^{5-3}=b ^{2}\]

OpenStudy (anonymous):

OHHH! i thought you put b^5-2. i was about to say 5-3=2

OpenStudy (anonymous):

okay i got it lol

OpenStudy (anonymous):

oh wait: my answer choices are different :( a. b^-8 b. b^8 c. 1 /b^-8 d. 1/b^8

OpenStudy (anonymous):

ohhhh okay i see what they did, i forgot about the add and subtracting rule. i put in my caculator -3-5 so it equaled -8

OpenStudy (anonymous):

so it must be b^-8? :)